# Differentially private projection-depth-based medians

### v2

@@ Abstract
We develop $(\epsilon,\delta)$-differentially private projection-depth-based medians using the propose-test-release (PTR) and exponential mechanisms. Under general conditions on the input parameters and the population measure, (e.g. we do not assume any moment bounds), we quantify the probability the test in PTR fails, as well as the cost of privacy via finite sample deviation bounds. We then present a new definition of the finite sample breakdown point which applies to a mechanism, and present a lower bound on the finite sample breakdown point of the projection-depth-based median. We demonstrate our main results on the canonical projection-depth-based median, as well as on projection-depth-based medians derived from trimmed estimators. In the Gaussian setting, we show that the resulting deviation bound matches the known lower bound for private Gaussian mean estimation. In the Cauchy setting, we show that the “outlier error amplification” effect resulting from the heavy tails outweighs the cost of privacy. This result is then verified via numerical simulations. Additionally, we present results on general PTR mechanisms and a uniform concentration result on the projected spacings of order statistics, which may be of general interest.
@@ 1 Introduction
Unlike the Tukey median and those studied by Ramsay et al., [21], when the location and scale measures are chosen appropriately, projection-depth-based medians simultaneously possess affine equivariance and achieve a high breakdown point [27, 28]. Affine equivariance is a critical property for location estimators, as it ensures that the units of measurement do not have an effect on the resulting estimate, see, for example, Zuo and Serfling, [29]. In addition, the breakdown point is a common measure of the robustness of an estimator. It measures the fraction of points which must be corrupted to make an estimate arbitrarily large [14]. For instance, the breakdown point of the Tukey median degrades quickly with the dimension, $(1+d)^{-1}$, whereas the projection-depth-based medians generally retain a much higher finite sample breakdown point, (e.g., approximately $1/2-d/n$ [27]). This finite sample breakdown point is carried over to the private setting, (see Theorem 3.3). One may also note that, when the location and scale measures are chosen appropriately, the projection-depth-based medians enjoy higher relative efficiency than that of the Tukey median [27]. Motivated by these properties, our work develops and studies differentially private projection-depth-based medians. We consider a general location estimation setting and do not place any moment bounds on the population measure.
@@
Our contributions are as follows: We first propose a class of $(\epsilon,\delta)$-differentially private projection-depth-based medians. Next, under general conditions on the population measure, we quantify both the probability that the PTR mechanism fails to release an estimate and the cost of privatizing the projection-depth-based medians (Theorems 3.1, and 3.2, respectively). To measure the cost of privacy, we provide finite sample bounds on the distance between a private projection-depth-based median and the corresponding population value, which hold with high probability. Next, we present a new definition of the finite sample breakdown point which applies to a mechanism, and present a lower bound on the finite sample breakdown point of the projection-depth-based median (Theorem 3.3). We then demonstrate our methods for the canonical projection-depth-based median, as well as on projection-depth-based medians derived from trimmed estimators. In particular, we show that Theorems 3.1 and 3.2 apply provided the “projected densities” are bounded below in an interval containing a range of inner quantiles (Condition 4). As a first example, we show that in the Gaussian setting, our deviation bound matches the optimal lower bound for sub-Gaussian mean estimation given by Kamath et al., [16] and Cai et al., [8] (Example 3.1). We then show that in the Cauchy setting, the increase in error resulting from contamination or heavy tails, which we call the “outlier error amplification” is larger than the cost of privacy (Example 3.2). We confirm this phenomenon with numerical studies (Section 4). Lastly, we provide some results for analysis of general PTR mechanisms (Section 5), as well as a uniform concentration result on the projected spacings of order statistics (Lemma C.2), which may be of general interest.
@@ 2.1 Outlyingness and projection-depth-based medians
We first review projection-depth-based medians, starting with projected outlyingness [9, 29, 27]. Let $\mathcal{M}_{1}(\mathbb{R}^{d})$ be the set of measures on $\mathscr{B}(\mathbb{R}^{d})$, where $\mathscr{B}(S)$ denotes the Borel sets of $S$. For a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ and $u$ from the unit $(d-1)$-sphere, $\mathbb{S}^{d-1}$, let $\nu_{u}$ be the law of $X^{\top}u$ if $X\sim\nu$.
@@
The projection-depth-based medians enjoy several nice properties, including local and global robustness, affine equivariance, strong $\sqrt{n}$-consistency, and a relatively simple limiting distribution [27, 28]. To the best of our knowledge, among the competing affine equivariant medians, when the location and scale measures are chosen appropriately, (e.g. the median and median absolute deviation), the projection-depth-based medians have the highest breakdown point [27]. We show later that the private projection-depth-based medians also admit high finite sample breakdown points, (Theorem 3.3). Projection-depth-based medians have been shown to have higher asymptotic efficiency, relative to competitors [27]. For more information on projection-depth-based medians, we refer the reader to [27, 28].
@@ 3.1 Private projection-depth-based medians
For brevity, we write $\operatorname{EM}_{\hat{\nu}}(\cdot;\tau,\eta,\epsilon)$ as simply $\operatorname{EM}_{\hat{\nu}}$ when $\tau,\eta,\epsilon$ are clear from the context. Now, this mechanism alone is not differentially private. That is because there are some datasets which are not “safe”. Safety, introduced by Brown et al., [6] for Gaussian mean estimation, is a localized version of the $(\epsilon,\delta)$-differential privacy guarantee. We present a definition specific to our context. (For a generalized definition, see Section 5.)
@@
$\operatorname{SM}(\hat{\nu};\epsilon,\delta)=\inf\{m\in\mathbb{Z}^{+}\colon\ \exists\tilde{\nu}\in\tilde{\mathcal{M}}(\hat{\nu},m)\ s.t.\ \tilde{\nu}\notin\mathbf{S}(\operatorname{EM}_{\hat{\nu}},\epsilon,\delta)\}.$
When $\epsilon$ and $\delta$ are clear from the context, we omit them, denoting the safety margin as $\operatorname{SM}(\hat{\nu}).$ The safety margin is the minimum number of points that we need to modify in the original dataset in order for it to become unsafe. Observe that if $\operatorname{SM}(\hat{\nu})>0$, then $\hat{\nu}$ is safe. Let $W$ be a standard Laplace random variable. The test is defined as follows:
@@
where if $Z=1$, then the test is passed.
@@
The private projection-depth-based median is $(\epsilon,\delta)$-differentially private (Theorem 5.1). Note that the mechanisms of Brown et al., [6] and Liu et al., 2021b [18] are PTR mechanisms based on exponential mechanisms and $Z$, using a different cost function. The performance of a PTR mechanism is naturally characterized by both the probability that the test is passed ($\mathbb{E}Z$) and the accuracy of the approximation of $\theta$ by $\tilde{\theta}$. With these two quantities, we characterize both the likelihood that the mechanism releases a statistic, and how useful that statistic is likely to be.
To analyze these quantities, we first introduce the projected sensitivity and the maximal absolute projected deviation of an arbitrary, univariate statistic, $T$, respectively defined as
@@
For all $d,n\geq 1$ there exists a sequence of functions non-increasing functions $\zeta_{n}\longrightarrow 0$ as $n\to\infty$ such that for all $t\geq 0$ and $T\in\{\mu,\sigma\}$,
$\Pr\Big{(}S_{n}(T)\geq t\Big{)}\lesssim\zeta_{n}(t).$
@@
$\mathbb{E}Z\geq 1-C\left(e^{\epsilon}\delta+a_{1}\exp\left(-\frac{a_{2}c_{2}n\eta}{4(\tau\vee 1)}\right)+\kappa_{n}+\zeta_{n}(c_{1}/4)\right).$
@@
Suppose Conditions 2 and 3 hold. Then, for all $n,d\geq 1$, $\epsilon,\eta>0$, $\delta\leq 1/2$, $\gamma~{}\in~{}(3\zeta_{n}(c_{1}/2\wedge c_{3})\vee 3(1-\mathbb{E}Z),1]$ and $\tau\geq 20(2c_{3}/(10c_{2}-c_{1})+c_{3}/c_{1})$ with probability $1-\gamma$, it holds that
@@
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{c_{2}\eta}{\epsilon}\left[\log(1/\gamma)\vee d\log\left(\frac{2c_{2}\tau}{c_{3}}+1\right)\vee\frac{c_{3}}{c_{1}}\right]\vee\frac{c_{2}}{c_{1}}\zeta_{n}^{-1}(\gamma/3).$
@@
The proof of Theorem 3.2 is given in Appendix B. Theorem 3.2 shows that, provided $\eta$ is relatively small and $\mathbb{E}Z$ is large enough, the cost of privacy for privately estimating the projection-depth-based median is governed by $\eta(\log(1/\gamma)\vee d)/\epsilon$. Together, Theorem 3.1 and Theorem 3.2 quantify the trade-off inherent in $\eta$. Smaller $\eta$ improves accuracy at the cost of a higher probability of failing the test. In general, it is desirable for the cost of privacy to be less than the sampling error, $c_{2}\zeta_{n}^{-1}(\gamma/3)/c_{1}$. We explore this in the following section.
@@
Before concluding this section, we present a result concerning the robustness of the private projection-depth-based median. We show that, if the test is passed, then the private projection-depth-based median has a finite sample breakdown point at least as large as that of $\mu$ and $\sigma$. Recall that the finite sample breakdown point of an estimator is the maximum fraction of points which can be set to arbitrary values, under which the estimator remains finite and bounded away from the boundary of the parameter space [10]. For a set $A$, let $\partial A$ denote the boundary of $A$ and if $A\subset\mathbb{R}^{d}$, for $y\in\mathbb{R}^{d}$, let $\lVert x-A\rVert=\inf_{x\in A}\lVert x-y\rVert$. Formally, for an estimator $T$ of parameter $\theta\in\Theta$,
@@
$\operatorname{BP}(T,\hat{\nu})=\sup\{\alpha\in[0,1/2)\colon\forall\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},\lfloor n\alpha\rfloor),\ \lVert T(\tilde{\nu})\rVert<\infty\text{ and}\ \lVert T(\tilde{\nu})-\partial\Theta\rVert>0\}.$
In the setting of differential privacy, the estimator is a random variable, and so we must modify the above definition. Let ^$\widehat{\mathcal{M}}_{1}(n,d)$ be the set of empirical measures constructed from datasets of size $n\times d$, $\{\nu_{\mathbb{X}_{n}}\colon\ \mathbb{X}_{n}\in\mathbf{D}_{n\times d}\}$.
Definition 3.4.
Given $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, the finite sample breakdown point of a mechanism $P$, which is used to estimate $\theta\in\Theta$ is given by
$\operatorname{BP}(P,\hat{\nu})=\sup\{\alpha\in[0,1/2)\colon\forall\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},\lfloor n\alpha\rfloor),\ \exists C>0\,\\ \ s.t.\ P_{\tilde{\nu}}(\{x\colon\lVert x\rVert<C,\ \lVert x-\partial\Theta\rVert>0\})=1\}.$
@@
The finite sample breakdown point of a mechanism is then the maximum fraction of points which can be set to arbitrary values, under which the estimator remains finite and bounded away from the boundary of the parameter space almost surely. We now present the following result, pertaining to the robustness of the projection-depth-based median.
@@
For all $0<\eta,\delta,\epsilon,\tau<\infty$ it holds that $\operatorname{BP}(\operatorname{EM}_{\hat{\nu}},\hat{\nu})\geq\operatorname{BP}(\mu,\hat{\nu})\wedge\operatorname{BP}(\sigma,\hat{\nu})$.
@@
We have $\Theta=\mathbb{R}^{d}$, and as such we need not consider the condition on $\partial\Theta$. Next, note that by definition $\operatorname{EM}_{\hat{\nu}}(\{x\colon O(x,\hat{\nu})\leq\tau\})=1$. Therefore, it suffices to show that the set $\{x\colon O(x,\tilde{\nu})\leq\tau\}$ is bounded. Consider $\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},k)$, where $k<n(\operatorname{BP}(\mu)\wedge\operatorname{BP}(\sigma))$. We have that
@@
Therefore, choosing high breakdown $\mu$ and $\sigma$ implies that the differentially private projection-depth-based median will also have high breakdown.

### v3

@@ Abstract
We develop $(\epsilon,\delta)$-differentially private projection-depth-based medians using the propose-test-release (PTR) and exponential mechanisms. Under general conditions on the input parameters and the population measure, (e.g. we do not assume any moment bounds), we quantify the probability the test in PTR fails, as well as the cost of privacy via finite sample deviation bounds. Next, we show that when some observations are contaminated, the private projection-depth-based median does not break down, provided its input location and scale estimators do not break down. We demonstrate our main results on the canonical projection-depth-based median, as well as on projection-depth-based medians derived from trimmed estimators. In the Gaussian setting, we show that the resulting deviation bound matches the known lower bound for private Gaussian mean estimation. In the Cauchy setting, we show that the “outlier error amplification” effect resulting from the heavy tails outweighs the cost of privacy. This result is then verified via numerical simulations. Additionally, we present results on general PTR mechanisms and a uniform concentration result on the projected spacings of order statistics, which may be of general interest.
@@ 1 Introduction
Unlike the Tukey median and those studied by Ramsay et al., [24], when the location and scale measures are chosen appropriately, projection-depth-based medians simultaneously possess affine equivariance and achieve a high breakdown point [31, 32]. Affine equivariance is a critical property for location estimators, as it ensures that the units of measurement do not have an effect on the resulting estimate, see, for example, Zuo and Serfling, [33]. In addition, the breakdown point is a common measure of the robustness of an estimator. It measures the fraction of points which must be corrupted to make an estimate arbitrarily large [16]. For instance, the breakdown point of the Tukey median lies between $(1+d)^{-1}$ and $1/3$ [21], whereas the projection-depth-based medians generally retain a much higher finite sample breakdown point, (e.g., approximately $1/2-d/n$ [31]). This finite sample breakdown point is carried over to the private setting, (see Theorem 3.4). One may also note that, when the location and scale measures are chosen appropriately, the projection-depth-based medians enjoy higher relative efficiency than that of the Tukey median [31]. In addition, [9] recently considered the behavior of a new subclass of projection-depth-based medians inspired by the median of means estimator. They demonstrate under general assumptions, e.g., the population measure need not have any moments, and under adversarial contamination this subclass of medians achieves sub-Gaussian rates with respect to Mahalanobis error. Motivated by these properties, our work develops and studies differentially private projection-depth-based medians. We consider a general location estimation setting and do not place any moment bounds on the population measure.
@@
Our contributions are as follows: We first propose a class of $(\epsilon,\delta)$-differentially private projection-depth-based medians. Next, under general conditions on the population measure, we quantify both the probability that the PTR mechanism fails to release an estimate and the cost of privatizing the projection-depth-based medians (Theorems 3.1, and 3.2, respectively). To measure the cost of privacy, we provide finite sample bounds on the distance between a private projection-depth-based median and the corresponding population value, which hold with high probability. Next, we show that when some observations are contaminated, the private projection-depth-based median does not break down, provided its input location and scale estimators do not break down (Theorem 3.4). We then demonstrate our methods for the canonical projection-depth-based median, as well as on projection-depth-based medians derived from trimmed estimators. In particular, we show that Theorems 3.1 and 3.2 apply provided the “projected densities” are bounded below in an interval containing a range of inner quantiles (Condition 4). As a first example, we show that in the Gaussian setting, our deviation bound matches the optimal lower bound for sub-Gaussian mean estimation given by Kamath et al., [18] and Cai et al., [8] (Example 3.1). We then show that in the Cauchy setting, the increase in error resulting from contamination or heavy tails, which we call the “outlier error amplification” is larger than the cost of privacy (Example 3.2). We confirm this phenomenon with numerical studies (Section 4). Lastly, we provide some results for analysis of general PTR mechanisms (Section 5), as well as a uniform concentration result on the projected spacings of order statistics (Lemma C.2), which may be of general interest.
@@ 2.1 Outlyingness and projection-depth-based medians
We first review projection-depth-based medians, starting with projected outlyingness [27, 10, 11, 33, 31]. Let $\mathcal{M}_{1}(\mathbb{R}^{d})$ be the set of measures on $\mathscr{B}(\mathbb{R}^{d})$, where $\mathscr{B}(S)$ denotes the Borel sets of $S$. For a measure $\nu\in\mathcal{M}_{1}(\mathbb{R}^{d})$ and $u$ from the unit $(d-1)$-sphere, $\mathbb{S}^{d-1}$, let $\nu_{u}$ be the law of $X^{\top}u$ if $X\sim\nu$.
@@
For brevity, we will refer to the projected outlyingness as simply the outlyingness, and denote $\mu(\nu_{u})$ as $\mu_{u}$, $\sigma(\nu_{u})$ as $\sigma_{u}$, $\mu(\hat{\nu}_{u})$ as $\hat{\mu}_{u}$, and $\sigma(\hat{\nu}_{u})$ as $\hat{\sigma}_{u}$. The sample version of the outlyingness function is given by $O(x,\hat{\nu})$. The canonical choices of $\mu$ and $\sigma$ are the median and the median absolute deviation, respectively. The outlyingness based on the median and the median absolute deviation was first developed by Stahel, [27] and Donoho, [10]. It was later studied in general by Zuo, [31]. Choosing $\mu$ and $\sigma$ to be the trimmed mean and trimmed variance recovers the objective function used in the PTR mechanism developed by Liu et al., 2021b [20] for differentially private mean estimation. Lastly, choosing $\mu$ and $\sigma$ to be the median of means and a median of means version of the mad recovers the outlyingness studied by Depersin and Lecué, [9]. In our examples, we consider the trimmed mean, the median, trimmed mean absolute deviation, median absolute deviation, and the interquartile range. The minimizer of the outlyingness function is a projection-depth-based median [27, 10, 31].
Definition 2.2.
Given a translation and scale equivariant estimator $\mu$ and a translation invariant and scale equivariant estimator $\sigma$, the projection-depth-based median of $\nu$ is
@@
where $P$ is uniform on $\operatorname*{{\rm argmin}}_{x\in\mathbb{R}^{d}}O(x,\nu)$.
The projection-depth-based medians enjoy several nice properties, including local and global robustness, affine equivariance, strong $\sqrt{n}$-consistency, and a relatively simple limiting distribution [31, 32]. To the best of our knowledge, among the competing affine equivariant medians, when the location and scale measures are chosen appropriately, (e.g. the median and median absolute deviation), the projection-depth-based medians have the highest breakdown point [31]. In Section 3, we show that when some observations are contaminated, the private projection-depth-based median does not break down, provided its input location and scale estimators do not break down (Theorem 3.4). Projection-depth-based medians have been shown to have higher asymptotic efficiency, relative to competitors [31]. In addition, recently, Depersin and Lecué, [9] showed that the median of means version of the projection-depth-based median exhibits sub-Gaussian rates, under adversarial contamination and fairly general assumptions. For more information on projection-depth-based medians, we refer the reader to [27, 10, 31, 32, 9].
@@ 3.1 Private projection-depth-based medians
For brevity, we write $\operatorname{EM}_{\hat{\nu}}(\cdot;\tau,\eta,\epsilon)$ as simply $\operatorname{EM}_{\hat{\nu}}$ when $\tau,\eta,\epsilon$ are clear from the context. Now, this mechanism alone is not differentially private. That is because there are some datasets which are not “safe”. Safety, introduced by Brown et al., [6] for Gaussian mean estimation, is a localized version of the $(\epsilon,\delta)$-differential privacy guarantee. We present a definition specific to our context. (For a generalized definition, see Section 5.) Take ^$\widehat{\mathcal{M}}_{1}(n,d)$ to be the set of empirical measures constructed from datasets of size $n\times d$, $\{\nu_{\mathbb{X}_{n}}\colon\ \mathbb{X}_{n}\in\mathbf{D}_{n\times d}\}$.
@@
$\operatorname{SM}(\hat{\nu})\coloneqq\operatorname{SM}(\hat{\nu};\epsilon,\delta)=\inf\{m\in\mathbb{Z}^{+}\colon\ \exists\tilde{\nu}\in\tilde{\mathcal{M}}(\hat{\nu},m)\ s.t.\ \tilde{\nu}\notin\mathbf{S}(\epsilon,\delta)\}.$
The safety margin is the minimum number of points that we need to modify in the original dataset in order for it to become unsafe. Observe that if $\operatorname{SM}(\hat{\nu})>0$, then $\hat{\nu}$ is safe. Let $W$ be a standard Laplace random variable. The test is defined as follows:
@@
where, if $Z=1$, then the test is passed.
@@
Note that if $\tilde{\theta}=\perp$, it is meant that nothing is output by the algorithm. That is, $\perp$ can be interpreted as a null value. The private projection-depth-based median is $(\epsilon,\delta)$-differentially private (Theorem 5.1). Note that the mechanisms of Brown et al., [6] and Liu et al., 2021b [20] are PTR mechanisms based on $Z$ and the exponential mechanism. Brown et al., [6] considers a different cost function and Liu et al., 2021b [20] consider several cost functions, one of which is the outlyingness based on the trimmed mean and variance. The performance of the proposed PTR algorithm is naturally summarized by both the probability that the test is passed ($\mathbb{E}Z$) and the accuracy of the approximation of $\theta$ by $\tilde{\theta}$. With these two quantities, we characterize both the likelihood that the mechanism releases a statistic, and how useful that statistic is likely to be.
To analyze these quantities, we first introduce the projected sensitivity and the maximal absolute projected deviation of an arbitrary, univariate estimator [^1] By univariate estimator, we mean a functional on the space $\mathcal{M}_{1}(\mathbb{R})$., $T$, respectively defined as
@@
For all $d,n\geq 1$ and $T\in\{\mu,\sigma\}$ there exists a sequence of non-increasing functions $\zeta_{n,T}\longrightarrow 0$ as $n\to\infty$ such that for all $t\geq 0$
$\Pr\Big{(}S_{n}(T)\geq t\Big{)}\lesssim\zeta_{n,T}(t).$
@@
$\mathbb{E}Z\geq 1-C\left(e^{\epsilon}\delta+a_{1}\exp\left(-\frac{a_{2}c_{2}n\eta}{4(\tau\vee 1)}\right)+\kappa_{n}+\zeta_{n,\sigma}(c_{1}/4)\right).$
@@
Suppose Conditions 2 and 3 hold. Then, for all $n,d\geq 1$, $\epsilon,\eta>0$, $\delta\leq 1/2$, $\gamma~{}\in~{}(3(\zeta_{n,\sigma}(c_{1}/2)\vee\zeta_{n,\mu}(c_{3}))\vee 3(1-\mathbb{E}Z),1]$ and $\tau\geq 20(2c_{3}/(10c_{2}-c_{1})+c_{3}/c_{1})$ with probability $1-\gamma$, it holds that
@@
$\lVert\tilde{\theta}-\theta\rVert\lesssim\frac{c_{2}\eta}{\epsilon}\left[\log(1/\gamma)\vee d\log\left(\frac{2c_{2}\tau}{c_{3}}+1\right)\vee\frac{c_{3}}{c_{1}}\right]\vee\frac{c_{2}}{c_{1}}\zeta_{n,\mu}^{-1}(\gamma/3).$
@@
The proof of Theorem 3.2 is given in Appendix B. Theorem 3.2 shows that, provided $\eta$ is relatively small and $\mathbb{E}Z$ is large enough, the cost of privacy for privately estimating the projection-depth-based median is governed by $\eta(\log(1/\gamma)\vee d)/\epsilon$. Together, Theorem 3.1 and Theorem 3.2 quantify the trade-off inherent in $\eta$. Smaller $\eta$ improves accuracy at the cost of a higher probability of failing the test. In general, it is desirable for the cost of privacy to be less than the sampling error, $c_{2}\zeta_{n,\mu}^{-1}(\gamma/3)/c_{1}$. We find this to be the case in the examples explored in the next section for robust choices of $\mu$ and $\sigma$. Indeed, for such choices, in the following examples, we have that $\kappa_{n}$ and $\zeta_{n,\sigma}(c_{1}/2)\vee\zeta_{n,\mu}(c_{3})$ are exponentially small in $n$, as well as $\eta$ can be taken to be $O(c_{d}/n)$, where $c_{d}$ is a term which is at most polynomial in the dimension, $d$. This results in a failure probability $\lesssim\delta e^{\epsilon}$ and a cost of privacy which is less than the sampling error.
Remark 3.3.
It is popular to consider the error of an estimator with respect to the Mahalanobis metric, e.g. [6, 9], as it accounts for the differing scales across different directions. However, we aim to develop a general theory which includes population measures for which the Mahalanobis metric may not be suitable. As a result, we develop results based on the Euclidean metric. However, Liu et al., 2021b [20] and Depersin and Lecué, [9] have considered the behavior of different subclasses of projection-depth-based medians (both privately and non-privately, respectively). Their results show good performance of these projection-depth-based medians, i.e., under their assumptions, they achieve sub-Gaussian rates. Furthermore, if we restrict to the median of means version of the projection-depth-based median and the assumptions of Depersin and Lecué, [9], one can straightforwardly substitute the techniques of Depersin and Lecué, [9] into the proof of Theorem 3.2 where appropriate, which would give the rate of convergence of the median of means version of the private projection-depth-based median under the Mahalanobis metric.
@@
Before concluding this section, we present a result concerning the robustness of the private projection-depth-based median. The finite sample breakdown point of an estimator is the maximum fraction of points which can be set to arbitrary values, under which the estimator lies in the interior of the parameter space [12]. For a set $A$, let $\operatorname{int}(A)$ denote the interior of $A$. Formally, for an estimator $T$ of parameter $\theta\in\Theta$, the finite sample breakdown point of $T$ with respect to $\hat{\nu}$ is
$\operatorname{BP}(T,\hat{\nu})=\sup\{\frac{m}{n}\colon\forall\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},m),\ T(\tilde{\nu})\in\operatorname{int}(\Theta)\}.$
Given $\hat{\nu}\in\widehat{\mathcal{M}}_{1}(n,d)$, define
@@
$\operatorname{BP}(\operatorname{EM},\hat{\nu})=\sup\{\frac{m}{n}\colon\forall\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},m),\exists\ C_{\hat{\nu},\alpha,\tau}<\infty\\ \text{such that}\operatorname{EM}_{\tilde{\nu}}(\{\lVert X\rVert<C_{\hat{\nu},\alpha,\tau}\})=1\}.$
@@
$\operatorname{BP}(\operatorname{EM},\hat{\nu})$ is the maximum fraction of points which can be set to arbitrary values, under which the estimate is bounded with probability 1. (Note that if $\lVert\tilde{\theta}\rVert<\infty$, then $\tilde{\theta}\in\operatorname{int}(\Theta)$.) We now present the following result, pertaining to the robustness of the private projection-depth-based median.
@@
For all $0<\eta,\delta,\epsilon,\tau<\infty$ it holds that $\operatorname{BP}(\operatorname{EM},\hat{\nu})\geq\operatorname{BP}(\mu,\hat{\nu})\wedge\operatorname{BP}(\sigma,\hat{\nu}).$
@@
Note that by definition $\operatorname{EM}_{\hat{\nu}}(\{x\colon O(x,\hat{\nu})\leq\tau\})=1$. Therefore, it suffices to show that the set $\{x\colon O(x,\tilde{\nu})\leq\tau\}$ is bounded for all $\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},k)$, where $k<n(\operatorname{BP}(\mu)\wedge\operatorname{BP}(\sigma))$. Consider $\tilde{\nu}\in\mathcal{M}_{1}(\hat{\nu},k)$, where $k<n(\operatorname{BP}(\mu)\wedge\operatorname{BP}(\sigma))$. We have that
@@
Therefore, if the data is contaminated such that $\mu$ and $\sigma$ do not break down, then Theorem 3.4 implies that the differentially private projection-depth-based median will also not breakdown.